If a + b = 43 and a + c = 30 and b + c = 55
a is the same number in all equations
b is the same number in all equations
c is the same number in all equations
what are a, b, and c
Thank you

Answer by checkley77(12569)

a+b=43 or a=43-b
a+c=30 or a=30-c now set these 2 equations
43-b=30-c
-b+c=30-43
-b+c=-13
b+c=55 now add these 2 equations
-----------------
2c=42
c=42/2
c=21 answer.
a=30-21
a=9 answer.
9=43-b
b=43-9
b=34 answer.
proofs:
b+c=55
34+21=55
a+b=43
9+34=43
43=43
a=c=30
9+21=30
30=30